OpenStudy (anonymous):

Which of the following is a true statement? m∠a = m∠c + m∠d , m∠a = m∠b + m∠d m∠a + m∠b = 90° , m∠a = m∠b + m∠c

OpenStudy (anonymous):



OpenStudy (unklerhaukus):

can you find m∠b?

OpenStudy (anonymous):

im not sure on how to do that.

OpenStudy (unklerhaukus):

you can find m∠b in terms of m∠c, m∠d, if you know the angle sum of a tringle , you can find m∠b in terms of angle m∠a if you know the angle sum of a straight line

OpenStudy (unklerhaukus):

*triangle

OpenStudy (anonymous):

360

OpenStudy (unklerhaukus):

the sum of the angles in a triangle is 180°,

OpenStudy (anonymous):

lol my bad -.- im always getting confused.

OpenStudy (unklerhaukus):

do you know the sum of angles on a straight line?

OpenStudy (anonymous):

isnt that 180?

OpenStudy (unklerhaukus):

yes!

OpenStudy (unklerhaukus):

so can you form two equations which use this informations?

OpenStudy (anonymous):

would you mind typing out the steps for me cause im still unsure on how to put it together?

OpenStudy (unklerhaukus):

using the angle sum of a triangle is 180°, we have m∠b + m∠c + m∠d =180°

OpenStudy (unklerhaukus):

can you form the other equation?

OpenStudy (anonymous):

Mmm not really..

OpenStudy (unklerhaukus):

m∠b + m∠a = . . .

OpenStudy (anonymous):

is it m<a+m<b=90?

OpenStudy (unklerhaukus):



OpenStudy (anonymous):

is that a yes?

OpenStudy (unklerhaukus):

its a no ,

OpenStudy (unklerhaukus):

90º is only half a straight line

OpenStudy (anonymous):

-.- i feel so dumb w. this question already..

OpenStudy (unklerhaukus):

m<a+m<b=

OpenStudy (anonymous):

my answer choices are m<a+m<b=90, m<a=m<c+m<d, m<a=m<b+m<d & m<a=m<b+m<c and the first and third are wrong..

OpenStudy (unklerhaukus):

m<a+m<b=180°

OpenStudy (unklerhaukus):

m∠b + m∠c + m∠d =180° solve this equation for m∠b by taking away m∠c + m∠d from both sides m<a+m<b=180° solve this equation for m∠b by taking m∠a from both sides

OpenStudy (unklerhaukus):

what do you get??

OpenStudy (anonymous):

ok where im confused is how is m<a+m<b=180?

OpenStudy (unklerhaukus):

all the way around is 360° half way around is 180°